OpenStudy (anonymous):

y = (e^x + 4^3x)^6 find the derivative

OpenStudy (anonymous):

\[y = (e^{x} + 4^{3x})^{6}\]

OpenStudy (anonymous):

(e\[is \it 6(e ^{x} + 4^{3x)^{5}} * (2e ^{x} + \ln(4)(3) + 4^{3x})\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[6(e^x+4^{3x})^5\times (e^x+4^{3x}\ln(4)\times 3)\]

OpenStudy (anonymous):

the derivative of \[e^x\] is \[e^x\]

OpenStudy (anonymous):

isnt it chain rule + product rule

OpenStudy (anonymous):

the derivative of \[b^x\] is \[b^x\times \ln(b)\] so by the chain rule the derivative of \[4^{3x}\] is \[4^{3x}\times \ln(4)\times 3\]

OpenStudy (anonymous):

i made a mistake i think take the product rule derivative of the inner

OpenStudy (anonymous):

no product here. i am fairly sure my answer is right. derivative is 6(blah)^5 times derivative of blah

OpenStudy (anonymous):

yep thx